1. Calculate the NNY if investment is 400, consumption change from 250 to 310 and NY from 7500 to 8500.
  2. Calculate the NNY if export is 4550, consumption was 350 and is now 475, taxes changed from 50 to 125 and NY was 5000 and is now 7000.
  3. Calculate NNY, if the investment is 375 consumption was 700 and is now 900, taxes changed from 70 to 175, imports changed from  100 to 175 and NY from 7500 to 9000.
  4. Income increase from O to 100 and consumption 60 to 130. What is the consumption function?
  5. Calculate the consumption if the income is O, 1500, 3000, and 4500, saving is -1000 and the savings change from 1000 to 1450.
  6. Calculate NNY, if the government spending is 300, consumption change from 150 to 275 and NY from 500-6000.
1
Expert's answer
2021-11-14T17:33:59-0500

Solution:

1.). NNY = (400 + 310 + 8500) – (400 + 250 + 7500) = 1,060

 

2.). NNY = (475 + 125 + 7000 +4550) – (350 + 50 + 5000 + 4550) = 2,200

 

3.). NNY = (375 + 900 + 175 – 175 + 9000) – (375 + 700 + 70 – 100 + 7500) = 1,730

 

4.). Consumption function:

Derive MPC = Change in income/ Change in consumption = 100 – 0/130 – 60 = 100/70 = 1.43

Consumption function: C = 60 + 1.43Y

 

5.). Income = Consumption + Saving

Consumption = Income – Saving

When income = 0

Consumption = 0 - -1000 = 1000

When income = 1500

Consumption = 1500 – 1000 = 500

When income = 3,000

Consumption = 3000 – 450 = 2550

When income = 4500

Consumption = 4500 – 1500 = 3000

6.).  NNY = (300 + 275 + 6000) – (300 + 150 + 500) = 5,625